An investment project has annual cash inflows of $4,900, $3,400, $4,600, and $3,800, for the next four years, respectively. The discount rate is 13 percent. intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Discounted payback period years What is the discounted payback period for these cash flows if the initial cost is $7,300? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Discounted payback period years What is the discounted payback period for these cash flows if the initial cost is $10,300? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Discounted payback period years
Answer #1

Discounted Payback Period =

( Last Year with a Negative Cumulative Cash Flow ) + [( Absolute Value of negative Cumulative Cash Flow in that year)/ Total Present Cash Flow in the following year)]

= 1+ (863.72 /2662.70)

= 1.32 Years

Hence the correct answer is 1.32 Years

Year Cash Flow Discounting Factor(13%) Present Value (Cash Flow * Discounting Factor) Cumulative Cash Flow
0 -5,200 1 -5200 -5,200
1 4,900 0.884955752 4,336.28 -863.72
2 3,400 0.783146683 2,662.70 1,798.98
3 4,600 0.693050162 3,188.03 4,987.01
4 3,800 0.613318728 2,330.61 7,317.62

At Inital Cost of $ 7,300

= 2 + (301.02 / $ 3,188.03)

= 2.09 Years

Year Cash Flow Discounting Factor(13%) Present Value (Cash Flow * Discounting Factor) Cumulative Cash Flow
0 -7,300 1 -7300 -7,300
1 4,900 0.884955752 4,336.28 -2,963.72
2 3,400 0.783146683 2,662.70 -301.02
3 4,600 0.693050162 3,188.03 2,887.01
4 3,800 0.613318728 2,330.61 5,217.62

at Initial Cost = $ 10,300

= 3 + ( 112.99/2330.61)

= 3.05 Years